OpenStudy (anonymous):

the nth no of a sequence is a(n)= 5^n + 12n - 1. Prove that it is divisible by 16 (using induction), considering a(n+1) - 5a(n).

OpenStudy (anonymous):

so i proved a(1), got a(n+1) and 5a(n). what do i do now??

OpenStudy (anonymous):

what is \(a_n\) ?

OpenStudy (anonymous):

5^n + 12n - 1.

OpenStudy (anonymous):

i just don't understand the last part of the q.

OpenStudy (anonymous):

i am sorry, i meant what is \(a_{n+1}\) ?

OpenStudy (anonymous):

btw, hi satellite :D

OpenStudy (anonymous):

hi! the idea is there will be some algebra involved to get \(a_{n+1}\) with part of it being \(a_n\) and the rest being 16 times something

OpenStudy (anonymous):

ohh :P umm, so, i'm saying that a(n) = 16m. so its 4(4m + 5^k + 3), with n = k, and substituting a couple of other things.

OpenStudy (anonymous):

let me see if i can do the algebra

OpenStudy (anonymous):

ohh, okay, go ahead :)

OpenStudy (anonymous):

ok it is like this:

OpenStudy (anonymous):

shoot.

OpenStudy (anonymous):

it is really a bunch of algebra, that is all you want to have a term that looks like \(5^n+12n-1\) when you compute \[5^{n+1}+12(n+1)-1\]so you force it in there

OpenStudy (anonymous):

here we go \[5^{n+1}+12n+12-1=5\times 5^n+12n+11\] \[=5(5^n+12n-1)+\text{stuff}\] so we just need to find out what the stuff is that makes it work now \[5(5^n+12n-1)=5^{n+1}+60n-5\] but you have \[5^{n+1}+12n+11\] so you have to subtract off \(56n\) and add \(16\) to make it match therefore you can say \[5^{n+1}+12(n+1)-1=5(5^n+12n-1)-56n+16\]

OpenStudy (anonymous):

my algebra is good, but my arithmetic is bad sorry it should be "subtract off \(48n\) and add 16

OpenStudy (anonymous):

so by this trick of algebra you get \[5^{n+1}+12(n+1)-1=5(5^n+12n-1)-48n+16\]

OpenStudy (anonymous):

therefore you have \[a_{n+1}=5(a_n)-16(3n-1)\] \(a_n\) is divisible by 16 "by induction" and \(16(3n-1)\) is divisible by 16 from your eyeballs, so the whole thing must be divisible by 16

OpenStudy (anonymous):

yup, i'm working it out.

OpenStudy (anonymous):

Thank youu SOO muchh. And it's good to see you (figuratively :P), cause i haven't been around in a while.